# Semi-tensor product-based convolutional neural network

### v2

@@ Abstract
keywords:
semi-tensor product, domain-based convolution, convolutional neural network, irregular image, higher-order signal identification
@@ 3.2 CP with finite windows
Note: although the same symbol “*” is often used in practice, (17) corresponds to cross-correlation rather than the commutative convolution in (15)-(16).
@@ 4.2 Block Hadamard product and receptive field matrix
Definition 4.3 (Block Hadamard Product).
@@
Definition 4.4 (Receptive Field Matrix (RFM)).
@@
Proof.
This follows directly from the construction of the RFM $R_{A}$ in (33) and the definition of the block Hadamard product in (30), which together implement the sliding-window inner product described in Algorithm 4.1. $\Box$
@@ 6.2 Partly damaged image
Example 6.2.
Consider the partly damaged image in Fig. 6, where the damaged region is indicated by $\times$”.
@@ 6.3 Proportional STP-CP
As mentioned in [22], “Assume the size of a dog is the same as the training dogs, CNN can recognize it well. But if the dog image is considerably enlarged, CNN may fail to recognize it.” Conventional CNNs often struggle with large-scale variations: a model trained on objects of one size may fail when objects are significantly larger or smaller. STP-CP addresses this by allowing receptive fields whose sizes do not match the kernel size exactly. By setting receptive field sizes appropriately, a proportional STP-CNN can recognize objects across scales. The following example illustrates proportional STP-CP for mismatched image sizes.
Example 6.3.
@@ 8 Conclusion
[…]
This work was supported by the National Natural Science Foundation of China (Grant No. 62350037).

### v3

@@ Abstract
Keywords semi-tensor product $\cdot$ domain-based convolution $\cdot$ convolutional neural network $\cdot$ irregular image $\cdot$ higher-order signal identification
@@ 3.2 CP with finite windows
Note: although the same symbol “*" is often used in practice, (17) corresponds to cross-correlation rather than the commutative convolution in (15)-(16).
@@ 4.2 Block Hadamard product and receptive field matrix
Definition 4.3 (Block Hadamard Product)
@@
Definition 4.4 (Receptive Field Matrix (RFM))
@@
Proof. This follows directly from the construction of the RFM $R_{A}$ in (33) and the definition of the block Hadamard product in (30), which together implement the sliding-window inner product described in Algorithm 4.1. $\Box$
@@ 6.2 Partly damaged image
Example 6.2
Consider the partly damaged image in Fig. 6, where the damaged region is indicated by $\times$".
@@ 6.3 Proportional STP-CP
As mentioned in [22], “Assume the size of a dog is the same as the training dogs, CNN can recognize it well. But if the dog image is considerably enlarged, CNN may fail to recognize it." Conventional CNNs often struggle with large-scale variations: a model trained on objects of one size may fail when objects are significantly larger or smaller. STP-CP addresses this by allowing receptive fields whose sizes do not match the kernel size exactly. By setting receptive field sizes appropriately, a proportional STP-CNN can recognize objects across scales. The following example illustrates proportional STP-CP for mismatched image sizes.
Example 6.3